# 3D-COCO: extension of MS-COCO dataset for scene understanding and 3D reconstruction

### v2

@@ 1 Introduction
At the same time, 3D reconstruction approaches based on neural networks have been developed. These architectures can be used, for example, in industrial or virtual reality applications. Recently, promising 3D reconstruction methods, such as 3D-C2FT [11], LegoFormer [12] or VPFusion [13]. Their performances are usually evaluated on ShapeNet [2]. Although this dataset includes a wide range of commonly encountered objects, it could be supplemented with new semantic classes present in detection datasets such as MS-COCO [1]. In addition, ShapeNet [2] only offers synthetic renderings, which limits the application of 3D reconstruction networks to real-world situations.
Thus, we propose 3D-COCO, an extension of the widely used MS-COCO [1] dataset, adapted for object detection configurable with text, 2D images, or 3D CAD model queries and for single or multi-view 3D reconstruction. The 3D-COCO dataset opens new perspectives to image detection by providing 3D models that are automatically aligned with 2D annotations. It also opens the way to the integration of real images for 3D reconstruction which until now remained confined to synthetic images. Moreover, 3D-COCO provides a wider variety of semantic classes for 3D reconstruction in addition to ShapeNet’s ones [2]. The Objaverse [3] 3D models database is used with ShapeNet [2] to provide a sufficient number of objects for each of the 80 MS-COCO [1] semantic classes. Alignment between collected 3D models and MS-COCO [1] 2D annotations is made using a simple, yet effective automatic class-driven retrieval method based on IoU.
To summarize, we make the following contributions:
@@
    Objaverse [3] * 80 N/A N/A $\approx$800K
    Objaverse XL [28] *** 80 N/A N/A $\approx$10, [^0] K
@@ 2 Related work
Among all these datasets, 3D models can be provided in many different formats :
  • Multi-view images in KITTI [18]
  • RGB-D images in SUN-RGBD [17]
  • Point clouds in Google Scans [24] and CO3D [25]
  • Meshes in IKEA [19], PASCAL3D+ [20], ObjectNet3D [21], ABO [22], 3DFuture [23], Pix3D [26] and PhotoShape [27]
  • Voxel grids in ShapeNet [2]
@@ 3.1 Collection of 3D models
The IoU-based matching method described in Figure 2 requires some pre-processing both on MS-COCO [1] annotations and on 3D-COCO models. Indeed, MS-COCO [1] annotations and API are used to generate a binary mask for each annotation as illustrated on the left part of Figure 2. This mask is saved on a $224\times 224$ image and normalized in scale to make the represented element touch the edges of the picture. For each CAD model, The same process has been made in 3.1 to get 62 binary masks represented on the bottom right side of Figure 1. This process allows to get silhouettes on images of similar sizes which will be compatible with an IoU calculation.
@@ 5 Conclusion
Nevertheless, for future iteration of the dataset, 3D-COCO could be improved with :
  • A better 2D-3D alignment method for articulated 3D models such as humans or animals.
  • Exploring other retrieval methods based on neural network feature extraction.
  • Integrating new 3D models to have a more balanced number of CAD models for each class.

### v3

@@ 1 Introduction
At the same time, 3D reconstruction approaches based on neural networks have been developed. These architectures can be used, for example, in industrial or virtual reality applications. Recently, promising 3D reconstruction methods, such as 3D-C2FT [11], LegoFormer [12] or VPFusion [13] hav emerged. Their performances are usually evaluated on ShapeNet [2]. Although this dataset includes a wide range of commonly encountered objects, it could be supplemented with new semantic classes present in detection datasets such as MS-COCO [1]. In addition, ShapeNet [2] only offers synthetic renderings, which limits the application of 3D reconstruction networks to real-world situations.
Thus, we propose 3D-COCO, an extension of the widely used MS-COCO [1] dataset, adapted for object detection configurable with text, 2D images, or 3D CAD (Computer-Aided Design) model queries and for single or multi-view 3D reconstruction. The 3D-COCO dataset opens new perspectives to image detection by providing 3D models that are automatically aligned with 2D annotations. It also opens the way to the integration of real images for 3D reconstruction which until now remained confined to synthetic images. Moreover, 3D-COCO provides a wider variety of semantic classes for 3D reconstruction in addition to ShapeNet’s ones [2]. The Objaverse [3] 3D models database is used with ShapeNet [2] to provide a sufficient number of objects for each of the 80 MS-COCO [1] semantic classes. Alignment between collected 3D models and MS-COCO [1] 2D annotations is made using a simple, yet effective automatic class-driven retrieval method based on IoU.
@@
Objaverse [3] N/A 80 N/A N/A $\approx$800K
Objaverse XL [28] N/A 80 N/A N/A $\approx$10, [^0] K
@@ 2 Related work
Among all these datasets, 3D models can be provided in many different formats : multi-view images in KITTI [18] ; RGB-D images in SUN-RGBD [17] ; point clouds in Google Scans [24] and CO3D [25] ; meshes in IKEA [19], PASCAL3D+ [20], ObjectNet3D [21], ABO [22], 3DFuture [23], Pix3D [26] and PhotoShape [27] ; voxel grids in ShapeNet [2].
@@ 3.1 Collection of 3D models
The IoU-based matching method described in Figure 2 requires some pre-processing both on MS-COCO [1] annotations and on 3D-COCO models. Indeed, MS-COCO [1] annotations and API are used to generate a binary mask for each annotation as illustrated on the left part of Figure 2. This mask is saved on a $224\times 224$ image and normalized in scale to make the represented element touch the edges of the picture. For each CAD model, we used the process described in 3.1 to get 62 binary masks represented on the bottom right side of Figure 1. This process allows to get silhouettes on images of similar sizes which will be compatible with an IoU calculation.
@@ 5 Conclusion
Nevertheless, for future iteration of the dataset, 3D-COCO could be improved with a better 2D-3D alignment method for articulated 3D models such as humans or animals. Exploring other retrieval methods based on neural network feature extraction or integrating new 3D models to have a more balanced number of CAD models for each class are also relevant perspectives.